Write the equation in standard from for the line that contains the given point with the given slope.
(1,0); m=3

Answer by darq(90)

(1,0); m=3

first use slope intercept point to find an equation.

y=mx+b
0=1*3+b
0=3+b
b=-3

so our equation will be

y=mx+b
y=3x-3

this is the slope intercept form.

y-3x=-3
-3x+y=-3
3x-y=3
now we have the standart form.